Solve the system of linear equations below.

6x + 3y = 33
4x + y = 15
A. x = 2, y = 7
B. x = -13, y = 7
C. x = - 2/3 , y = 12 2/3
D. x = 5, y = 1

Answer by ewatrrr(24785)

 
Hi
6x + 3y = 33
4x + y = 15 OR y = (-4x+15)
Solving by substitution:
6x + 3(-4x+15) = 33
12 = 6x
2 = x and y = 7 (-4·2 +15)